OpenStudy (anonymous):

WILL FAN AND MEDAL!!~~ Solve each equation for the indicated variable: 2x+y=8;y How do I solve this?!?!?

OpenStudy (anonymous):

What are the equations?

OpenStudy (anonymous):

In other words, it wants you to put 'y' at one side of the equal sign, and the rest at the other side.

OpenStudy (anonymous):

2x + y = 8 It wants you to have 'y' on one side and have '2x' and '8' on another side. So to get rid of 2x on the side 'y' is on, you would subtract it because '2x' is positive, so you will do the opposite. You will end up with y = 8 - 2x That's how you solve for 'y'.

OpenStudy (anonymous):

whats your number?

OpenStudy (anonymous):

I dont get it.... @greenlegodude57

OpenStudy (anonymous):

@austin_coleman

OpenStudy (anonymous):

What are your equations?

OpenStudy (anonymous):

2x+y=8;y

OpenStudy (anonymous):



OpenStudy (anonymous):

equation is 2x+y=8 solve for y

OpenStudy (anonymous):

I already gave you the answer..you just have to subtract 2x..

OpenStudy (anonymous):

y = 8 - 2x is your answer.

OpenStudy (anonymous):

There is no single number answer, because you don't know what x or y is.

OpenStudy (anonymous):

gotcha thanks

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

@Mychelynn99 .... @greenlegodude57 is correct

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

:) no prob